OpenStudy (anonymous):

Find the linear approximation L(x) of ln(x) at the point a = 12?

OpenStudy (anonymous):

I got 1/12 when I found the derivative and plugged in, but I know it's not right

OpenStudy (loser66):

at the point a =12, you have f(a) =ln (12). and f'(x) = 1/x so f'(8) = 1/8 plug them onto the formula \[\huge L(a) = f(a) + f'(a)*(x-a)\] you have the formula to calculate some value of function near 8.